OpenStudy (anonymous):

The circle is centered at the origin, and the length of its radius is 6. What is the circle's equation?

OpenStudy (raden):

use the formula ; x^2 + y^2 = r^2 here given r = 6

OpenStudy (anonymous):

is it x2 + y2 = 36?

OpenStudy (raden):

yes, that's right

OpenStudy (anonymous):

Thank you so much(: